y=3x+1; y=5x-3x how do I substitute
Answer by stanbon(75887)

y=3x+1; y=5x-3 how do I substitute
-----
Both expressions are "y", so equate them:
3x+1 = 5x-3
2x = 4
x = 2
================
Since y = 3x+1
y = 3*2+1
y = 7
===========
Cheers,
Stan H.